OpenStudy (anonymous):

y + 2x = −1 3y − x = 4 graph of two lines that intersect at the point (negative 1, 1) graph of two lines that intersect at the point (1, 1) graph of two parallel lines with positive slopes None of the above

OpenStudy (anonymous):

First, we need to put this is slope intercept form. that is y=mx+b. Do do that, take the first equation: y+2x=-1 subtract -2x from both sides You're left with y=-2x-1 Now do you now how to do the same for the second equation?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

To put it in slope intercept for, move the x to the other side, it'll now be: 3y=x+4 Now divide the equation by 3, and you get: y=1/3x+4/3 So you're two equations are : y=-2x-2 y=1/3x+4/3 When you graph them , they intersect at (-1,1). So it is a.

OpenStudy (anonymous):

I meant y=-2x-1 for the first equation...typo XD

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome